Tom's age is thrice as Alice's age.If the sum of their ages is 54, What is the age of tom and Alice?
Answer by Schaman_Dempster(26)

Let the ages of Tom and Alice be x and y respectively.
We know that Tom's age is thrice as Alice's age.
So, x = 3y --------------- (1)
Also, we know that the sum of their ages is 54.
Therefore,
x + y = 54
or 3y + y = 54 { Using equation (1) }
4y = 54
or y = 54/4 = 13.5
Plugging back the value of y in equation (1), we get:
x = 3*13.5 = 40.5
Hence, Tom's age is 40.5 years and Alice's age is 13.5 years
Hope this helps~!